On January 1, Year 1, Price, Co. issued $368,000 of five-year, 6 percent bonds at 97. Interest payable is annually on December 31. The discount is amortized using the straight-line method.

Prepare journal entries to record the bond transactions for Year 1 and 2, with year 1 on this table and year two on another. If no entry is required for a transaction/event, select "no journal entry required". Note: enter debits before credits.

Date General Journal Debit Credit
Jan 01
Answer #1
Date Debit Credit
Jan 01 Cash 356960 =368000*0.97
Discount on Bonds payable 11040
       Bonds payable 368000
Dec 31 Interest expense 24288
     Discount on Bonds payable 2208 =11040/5
     Cash 22080 =368000*6%
Year 2
Date Debit Credit
Dec 31 Interest expense 24288
     Discount on Bonds payable 2208 =11040/5
     Cash 22080 =368000*6%